# MaDiS: Taming Masked Diffusion Language Models for Sign Language Generation

### v1

@@ 3.3 Supervised Fine-Tuning
Figure 3: (a) Vanilla MDLMs adopt an unconstrained confidence-based unmasking strategy [45]. (b) We insert temporal checkpoints at noise levels $0.75$ and $0.5$ to reduce the complexity of unmasking orders. (c) The original unmasking strategy allows arbitrary unmasking orders (about $10^{123}$ for generating 100 tokens), while our method reduces this complexity by over $10^{41}$ times.
@@ 3.3.1 Unmasking with Temporal Checkpoints (UTC)
During the reverse process, the original MDLM [45] adopts a simple confidence-based unmasking strategy (Figure 3a). Specifically, at each step, the model unmasks $k$ tokens with the highest confidence, defined as the maximum of their categorical probabilities. However, this unconstrained strategy leads to highly diverse unmasking orders. Denoting the number of generated tokens as $M$, the total number of possible unmasking orders is:
$N_{u}=\binom{M}{k}\binom{M-k}{k}\cdots\binom{k}{k}=\frac{M!}{(k!)^{M/k}}.$ (6)
@@
When $M=100$ and $k=4$, this number is approximately $2.9\times 10^{123}$. Although such diversity can enhance model performance, it also results in slow convergence [52].
@@
Inspired by VAR [70], which adopts a coarse-to-fine generation scheme with progressively enhanced resolution, we propose a novel unmasking strategy with temporal checkpoints (UTC), as shown in Figure 3b. Specifically, we divide the reverse process into three stages at $t=0.75$ and $t=0.5$, where $1/4$ and $1/2$ of the tokens have been unmasked, respectively. Both pivot states are enforced to contain uniformly distributed tokens, while a standard confidence-based unmasking strategy is applied within each stage. Consequently, the number of possible orders under our UTC is substantially reduced to:
$N_{u}^{\mathrm{UTC}}=\frac{(M/4)!}{(k!)^{M/4k}}\cdot\frac{(M/4)!}{(k!)^{M/4k}}\cdot\frac{(M/2)!}{(k!)^{M/2k}}.$ (7)
@@
When $M=100$ and $k=4$, UTC reduces the number of orders by $10^{41}\times$ compared to vanilla unmasking strategy (Figure 3c), thereby improving training efficiency. Moreover, the model can generate coarse-grained results at lower temporal resolutions in the early stages of inference, which provides a basis for further acceleration. To ensure training–inference consistency, we employ predefined index masks to filter out infeasible states during training. More implementation details are provided in the supplement.
3.3.2 Mixture-of-Parts (MoP) Embedding Layer
@@
Figure 4: (a) A typical sign embedding layer averages part-wise token embeddings [96]. (b) The proposed mixture-of-parts embedding layer leverages optimized VAE codebooks and uses a learnable gate to control contributions from different body parts.
@@
In contrast to the pretraining stage, where different part-wise tokens are treated equally, we aim to more discriminatively capture the relations among body parts during fine-tuning with text conditions. To this end, our MoP embedding layer is built upon the well-trained VAE codebooks and employs a learnable gate to control the contributions from different parts. For instance, since numerous signs are produced with a single hand [7, 22], it would be beneficial for the model to assign distinct weights to each hand.
@@
As shown in Figure 4b, given the part IDs ($i_{p}$), the part-wise code embeddings are passed through fully-connected (FC) layers to match the model dimension of MDLM: $\mathbf{E}_{p}=\mathrm{FC}_{p}(\mathbf{c}_{p}^{i_{p}})$. Furthermore, we use an MLP with a softmax layer to learn gating weights: $\mathbf{G}=\{g_{p}\}=\mathrm{softmax}(\mathrm{MLP}(\mathrm{cat}(\{\mathbf{c}_{p}^{i_{p}}\})))$. The final sign embedding is obtained as $\mathbf{E}=\sum_{p}g_{p}\cdot\mathbf{E}_{p}$.
@@
Similar to [34], we also observe that continuing to use pretraining objectives during fine-tuning is beneficial. Therefore, the fine-tuning loss is defined as $\mathcal{L}_{sft}=\mathcal{L}_{tok}+\alpha(\mathcal{L}_{lat}+\mathcal{L}_{phy})$, where $\alpha=0.5$ by default.
@@ 3.4 SiBLEU and SiCLIP
Existing SLG models [96, 5, 97, 1] are typically evaluated using dynamic time warping (DTW) error [1] and back-translation (BT) [59]. However, DTW error may over-optimistically reflect model performance since it measures the minimum distance between generated and ground-truth sequences. Moreover, its computation is inefficient due to the reliance on dynamic programming. On the other hand, BT evaluates performance in the text space, which overlooks the visual and multi-cue nature of sign language, and the translation process itself introduces unavoidable errors [30]. To address these issues, we propose two new metrics that serve as effective alternatives.
SiBLEU – Straightforward and Efficient Evaluation. For LM-based SLG approaches, signs are tokenized, which enables a straightforward evaluation by computing BLEU scores [48] over generated tokens against ground-truth tokens. We term this the SiBLEU score. A higher SiBLEU indicates greater overlap with the ground truth.
SiCLIP – Evaluation in a Joint Embedding Space. In related domains such as human motion generation [69, 42] and text-to-video generation [89, 78, 80], CLIP [55] is widely used to evaluate the alignment between prompts (texts) and generations in a joint-embedding space. Motivated by this, we train a CLIP model over ground-truth sign-text pairs, and adopt retrieval-based metrics [8, 50, 33] for evaluating the semantic alignment between texts and generated signs. More details are provided in the supplement.
Figure 5: Qualitative comparisons of generated signs between our proposed method, MaDiS, with the SOTA method, SOKE [96], on the test sets of CSL-Daily (left), Phoenix-2014T (middle), and How2Sign (right).
@@ 4 Experiments
Datasets. We evaluate MaDiS on three widely-adopted datasets: CSL-Daily [94], Phoenix-2014T [11], and How2Sign [18], which contain 20K/8K/35K video-text pairs for Chinese, American, and German sign languages, respectively. Since we focus on generating sign motions, we use the curated SMPL-X poses provided in [96].
Evaluation Metrics. Due to the length difference between generated signs and ground truth, we follow the previous work [96] to use dynamic time warping over joint position errors (DTW-JPE) as the major metric to measure sequence-level distances. Note that we do not compute procrustes alignment since it will align rotation before computing joint errors, leading to overoptimistic performance [9]. As a complement, we use SiBLEU-4 as an indicator for token-level precision and report the sentence-level sign-to-text retrieval performance (R@1) [8, 16] using SiCLIP as a joint feature extractor for sign motions and texts.
Implementation Details. We instantiate the MDLM with Qwen3-0.6B-Base [81], a leading open-source LLM supporting 119 languages. To enable bidirectional modeling, we replace the causal attention mask in each transformer layer with a non-causal one [45]. During pretraining, the model is trained on the combined training sets of all three datasets to maximize data utilization [29]. In the fine-tuning stage, it is trained on each dataset independently. We empirically set $k=4$ and $M=100$, which corresponds to 400 frames after decoding due to the 4x up-sampling layer in the VAE decoder. This length is sufficient to cover about 98% of training videos, and over-length videos are uniformly sampled to fit within 400 frames. When generating sign tokens, we adopt multi-head decoding [96] to predict all three part-wise tokens simultaneously, and tokens appearing after the end-of-sentence token are discarded [45]. The initial learning rate is set to 2e42e-4, and we adopt the AdamW optimizer [39] with a cosine learning rate scheduler and a batch size of 64 per GPU for both stages. The models are trained for 200 and 150 epochs in the pretraining and fine-tuning stages, respectively, using 4x Nvidia GH200 GPUs.
@@
Prog. Trans. [59] 16.30 32.63 N/A 13.65 15.01 31.77 N/A 14.03 14.74 30.17 N/A 9.88
Text2Mesh [66] 13.76 30.37 N/A 15.84 14.04 31.64 N/A 12.08 15.50 32.97 N/A 9.30
T2S-GPT [84] 12.32 15.43 2.11 N/A 20.03 11.65 19.09 1.93 N/A 22.10 12.65 18.44 2.13 N/A 12.58
NSA [5] N/A N/A 9.16 18.51 N/A 15.04
S-MotionGPT [32] 11.58 11.31 2.26 N/A 20.37 10.42 9.08 1.97 N/A 32.42 12.41 13.74 2.32 N/A 12.77
MoMask++ [26] 8.93 10.57 1.53 N/A 22.62 8.16 9.96 1.49 N/A 37.88 9.06 11.34 1.58 N/A 20.39
SOKE (mBART) [96] 7.38 9.68 2.59 2.14 23.64 6.04 7.72 2.71 1.85 34.30 7.75 10.08 2.52 1.74 16.12
SOKE (Qwen3) [96] 7.29 9.59 2.71 2.27 25.05 6.09 7.65 2.38 1.98 38.53 7.18 10.09 2.57 1.97 18.48
@@
Table 1: Comparison with SOTA sign language generation methods. Note that since SiBLEU scores are computed over sign tokens, we report “N/A” for tokenizer-free methods [59, 66, 5]. Similarly, for methods [84, 32, 26] that employ a single tokenizer for whole-body motions, the SiBLEU-hands results are also marked as “N/A”. DTW errors are reported by [96]. Methods using external sign dictionaries.
@@ 4.1 Comparison with State-of-the-Art Methods
Qualitative Comparison. We first conduct qualitative comparisons with the previous SOTA method, SOKE [96], which proposes a simple autoregressive SLG model without dedicated pretraining and can only capture unidirectional contexts due to the nature of ARLMs. In contrast, our MaDiS is pretrained across multiple sign representation spaces and benefits from an MDLM backbone that enables bidirectional contextual modeling. As shown in Figure 5, our model produces more precise and natural sign motions, demonstrating better generalization across diverse hand shapes, palm orientations, single- and double-handed signs, and finger details.
Method CSL-Daily Phoenix-2014T How2Sign
SOKE [96] 1.074 1.019 1.021
MDLM baseline 0.784 0.723 0.735
MaDiS (ours) 0.782 0.724 0.734
Table 2: Comparison of inference latency (s/video) with SOKE. The generated sequence length is fixed at 100 tokens. All measurements are conducted on a single GH200 GPU.
@@
Quantitative Comparison. As detailed in Table 1, we conduct a quantitative comparison with SOTA SLG methods. Since we introduce two new evaluation metrics, we carefully reproduce previous works using their open-source codes [59, 66, 32, 26, 96] or our own reimplementations [84, 5]. MoMask++ [26] is a leading human motion generation method based on a masked generation approach [13]. However, its masking schedule is heuristically determined without a theoretical foundation, and its multi-scale tokenization introduces error accumulation during inference, leading to inferior performance than SOKE. For fairness, we also report the performance of SOKE when using Qwen3-0.6B as the LM, instead of the originally adopted mBART-Large [38]. We observe that employing Qwen3 leads to slightly better performance, which can be attributed to the larger pretraining corpus of the Qwen3 series. Furthermore, we report the performance of an MDLM baseline, obtained by directly training an MDLM on sign language datasets. Notably, even without any pretraining techniques, this baseline performs on par with SOKE, which utilizes external sign dictionaries, while offering reduced latency (Table 2). This highlights the potential of masked diffusion language models for SLG. Finally, our proposed MaDiS achieves new SOTA results across the test sets of all three datasets and all evaluation metrics.
Efficiency. Real-world applications typically demand low-latency systems. We evaluate the inference latency of our model against previous ARLM-based SLG method, SOKE, as reported in Table 2. Across the three datasets, our method consistently achieves around 30% reduction in latency, attributable to the highly efficient parallel token generation enabled by the reverse process in MDLMs.

### v2

@@ 3.3 Supervised Fine-Tuning
Figure 3: (a) Vanilla MDLMs adopt an unconstrained confidence-based unmasking strategy [llada]. (b) We insert temporal checkpoints at noise levels $0.75$ and $0.5$ to reduce the complexity of unmasking orders. (c) The original unmasking strategy allows arbitrary unmasking orders (about $10^{123}$ for generating 100 tokens), while our method reduces this complexity by over $10^{41}$ times.
@@
Unmasking with Temporal Checkpoints (UTC). During the reverse process, the original MDLM [llada] adopts a simple confidence-based unmasking strategy (Fig.˜3a). Specifically, at each step, the model unmasks $k$ tokens with the highest confidence, defined as the maximum of their categorical probabilities. However, this unconstrained strategy leads to highly diverse unmasking orders. Denoting the number of generated tokens as $M$, the total number of possible unmasking orders is:
$N_{u}=\binom{M}{k}\binom{M-k}{k}\cdots\binom{k}{k}=\frac{M!}{(k!)^{M/k}}.\vskip-5.69054pt$ (4)
@@
When $M=100$ and $k=4$, this number can achieve $2.9\times 10^{123}$. Although such diversity can enhance model performance, it also results in slow convergence [prabhudesai2025diffusion].
@@
Inspired by VAR [tian2024var], which adopts a coarse-to-fine generation scheme with progressively enhanced resolution, we propose a novel unmasking strategy with temporal checkpoints (UTC), as shown in Fig.˜3b. Specifically, we divide the reverse process into three stages at $t=0.75$ and $t=0.5$, where $1/4$ and $1/2$ of the tokens have been unmasked, respectively. Both pivot states are enforced to contain uniformly distributed tokens, while a standard confidence-based unmasking strategy is applied within each stage. Consequently, the number of possible orders under our UTC is substantially reduced to:
$N_{u}^{\mathrm{UTC}}=\frac{(M/4)!}{(k!)^{M/4k}}\cdot\frac{(M/4)!}{(k!)^{M/4k}}\cdot\frac{(M/2)!}{(k!)^{M/2k}}.\vskip-2.84526pt$ (5)
@@
When $M=100$ and $k=4$, UTC reduces the number of orders by $10^{41}\times$ compared to vanilla unmasking strategy (Fig.˜3c). Moreover, the resulting generation process follows a coarse-to-fine structure, with both factors jointly easing model learning and accelerating convergence. To ensure training-inference consistency, we employ predefined index masks to filter out infeasible states during training. More implementation details are provided in the supplement.
@@
Figure 4: (a) A typical sign embedding layer averages part-wise token embeddings [zuo2025soke]. (b) The proposed mixture-of-parts embedding layer leverages optimized VAE codebooks and uses a learnable gate to control contributions from different body parts.
@@
Mixture-of-Parts (MoP) Embedding Layer. In contrast to the pretraining stage, where different part-wise tokens are treated equally, we aim to more discriminatively capture the relations among body parts during fine-tuning with text conditions. To this end, our MoP embedding layer is built upon the well-trained VAE codebooks and employs a learnable gate to control the contributions from different parts. For instance, since numerous signs are produced with a single hand [battison1978lexical, forte2023reconstructing], it would be beneficial to assign distinct weights to each hand.
@@
As shown in Fig.˜4b, given the part IDs ($i_{p}$), the part-wise code embeddings are passed through fully-connected (FC) layers to match the model dimension of MDLM: $\mathbf{E}_{p}=\mathrm{FC}_{p}(\mathbf{c}_{p}^{i_{p}})$. Furthermore, we use an MLP with a softmax layer to learn gating weights: $\mathbf{G}=\{g_{p}\}=\mathrm{softmax}(\mathrm{MLP}(\mathrm{cat}(\{\mathbf{c}_{p}^{i_{p}}\})))$. The final sign embedding is obtained as $\mathbf{E}=\sum_{p}g_{p}\cdot\mathbf{E}_{p}$.
@@
Similar to [jiao2024visual], we also observe that continuing to use pretraining objectives during fine-tuning is beneficial. Therefore, the fine-tuning loss is defined as $\mathcal{L}_{sft}=\mathcal{L}_{tok}+\alpha(\mathcal{L}_{lat}+\mathcal{L}_{phy})$, where $\alpha=0.5$ is determined empirically through a sensitivity analysis in the supplement.
@@ 3.4 SiBLEU and SiCLIP
Existing SLG models [zuo2025soke, baltatzis2024nsa, zuo2024simple, arkushin2023ham2pose] are typically evaluated using dynamic time warping (DTW) error [arkushin2023ham2pose] and back-translation (BT) [saunders2020progressive]. However, DTW error may over-optimistically reflect model performance since it measures the minimum distance between generated and ground-truth sequences. Moreover, its computation is inefficient due to the reliance on dynamic programming. On the other hand, BT evaluates performance in the text space, which overlooks the visual and multi-cue nature of sign language, and the translation process itself introduces unavoidable errors [SiLVERScore]. To address these issues, we propose two new metrics that serve as effective alternatives.
SiBLEU – Straightforward and Efficient Evaluation in the Token Space. For LM-based SLG approaches, signs are tokenized, which enables a straightforward evaluation by computing BLEU scores [papineni2002bleu] over generated tokens against ground-truth tokens. We term this the SiBLEU score. A higher SiBLEU indicates greater overlap with the ground-truth tokens.
SiCLIP – Evaluation in a Joint Embedding Space. In related domains such as human motion generation [tevet2022motionclip, meng2025rethinking] and text-to-video generation [zhang2024controlvideo, wu2023tune-a-video, xing2024simda], CLIP [clip] is widely used to evaluate the alignment between prompts (texts) and generations in a joint-embedding space. Motivated by this, we train a CLIP model over ground-truth sign-text pairs in the training sets, and adopt retrieval-based metrics [bensabath2025text, petrovich2023tmr, jiang2024signclip] for evaluating the semantic alignment between texts and generated signs. More details are provided in the supplement.
Figure 5: Qualitative comparisons of generated signs between our proposed method, MaDiS, with the SOTA method, SOKE [zuo2025soke], on the test sets of CSL-Daily (left), Phoenix-2014T (middle), and How2Sign (right).
@@ 4 Experiments
Datasets. We evaluate MaDiS on three widely-adopted datasets: CSL-Daily [csl-daily], Phoenix-2014T [2014T], and How2Sign [duarte2021how2sign], which contain 20K/8K/35K video-text pairs for Chinese, American, and German sign languages, respectively. We use the curated SMPL-X poses provided in [zuo2025soke] as sign motion representations.
Evaluation Metrics. Due to the length difference between generated signs and ground truth, we follow the previous work [zuo2025soke] to use dynamic time warping over joint position errors (DTW-JPE) as the major metric to measure sequence-level distances. Note that we do not compute procrustes alignment (PA) since it will align rotation before computing joint errors, leading to overoptimistic performance [black2023bedlam]. As a complement, we use SiBLEU-4 as an indicator for token-level precision and report the sentence-level sign-to-text retrieval performance (R@1) [bensabath2025text, cheng2023cico] using SiCLIP as a joint feature extractor for signs and texts.
Implementation Details. We instantiate the MDLM with Qwen3-0.6B-Base [yang2025qwen3], a leading open-source LLM supporting 119 languages. To enable bidirectional modeling, we replace the causal attention mask in each transformer layer with a non-causal one [llada]. During pretraining, the model is trained on the combined training sets of all three datasets to maximize data utilization [hu2023signbert+, zhang2024scaling]. In the fine-tuning stage, it is trained on each dataset independently. We empirically set $k=4$ and $M=100$, which corresponds to 400 frames after decoding due to the 4x up-sampling layer in the VAE decoder. This length is sufficient to cover about 98% of training videos, and over-length videos are uniformly sampled to fit within 400 frames. When generating sign tokens, we adopt multi-head decoding [zuo2025soke] to predict all three part-wise tokens simultaneously, and tokens appearing after the end-of-sentence token are discarded [llada]. The initial learning rate is set to 2e42e-4, and we adopt the AdamW optimizer [adamw] with a cosine learning rate scheduler and a batch size of 64 per GPU for both stages. The models are trained for 200 and 150 epochs in the pretraining and fine-tuning stages, respectively, using 4x Nvidia GH200 GPUs.
@@
Prog. Trans. [saunders2020progressive] 16.30 32.63 N/A 13.65 15.01 31.77 N/A 14.03 14.74 30.17 N/A 9.88
Text2Mesh [stoll2022there] 13.76 30.37 N/A 15.84 14.04 31.64 N/A 12.08 15.50 32.97 N/A 9.30
T2S-GPT [yin-etal-2024-t2s] 12.32 15.43 2.11 N/A 20.03 11.65 19.09 1.93 N/A 22.10 12.65 18.44 2.13 N/A 12.58
NSA [baltatzis2024nsa] N/A N/A 9.16 18.51 N/A 15.04
S-MotionGPT [jiang2023motiongpt] 11.58 11.31 2.26 N/A 20.37 10.42 9.08 1.97 N/A 32.42 12.41 13.74 2.32 N/A 12.77
MoMask++ [guo2025snapmogen] 8.93 10.57 1.53 N/A 22.62 8.16 9.96 1.49 N/A 37.88 9.06 11.34 1.58 N/A 20.39
SOKE (mBART) [zuo2025soke] 7.38 9.68 2.59 2.14 23.64 6.04 7.72 2.71 1.85 34.30 7.75 10.08 2.52 1.74 16.12
SOKE (Qwen3) [zuo2025soke] 7.29 9.59 2.71 2.27 25.05 6.09 7.65 2.38 1.98 38.53 7.18 10.09 2.57 1.97 18.48
@@
Table 1: Comparison with SOTA sign language generation methods. Note that since SiBLEU scores are computed over sign tokens, we report “N/A” for tokenizer-free methods [saunders2020progressive, stoll2022there, baltatzis2024nsa]. Similarly, for methods [yin-etal-2024-t2s, jiang2023motiongpt, guo2025snapmogen] that employ a single tokenizer for whole-body motions, the SiBLEU-Hands results are also marked as “N/A”. DTW errors are reported by [zuo2025soke]. Methods using external sign dictionaries.
@@ 4.1 Comparison with State-of-the-Art Methods
Qualitative Comparison. We first conduct qualitative comparisons with the previous SOTA method, SOKE [zuo2025soke], which proposes a simple autoregressive SLG model without dedicated pretraining and can only capture unidirectional contexts due to the nature of ARLMs. In contrast, our MaDiS is pretrained across multiple sign representation spaces and benefits from an MDLM backbone that enables bidirectional contextual modeling. As shown in Fig.˜5, our model produces more precise and natural sign motions, demonstrating better generalization across diverse hand shapes, palm orientations, single- and double-handed signs, and finger details. Video demos are provided in the supplement.
@@
Quantitative Comparison. As detailed in Tab.˜1, we conduct a quantitative comparison with SOTA SLG methods. Since we introduce two new evaluation metrics, we carefully reproduce previous works using their open-source codes [saunders2020progressive, stoll2022there, jiang2023motiongpt, guo2025snapmogen, zuo2025soke] or our own reimplementations [yin-etal-2024-t2s, baltatzis2024nsa]. MoMask++ [guo2025snapmogen] is a leading human motion generation method based on a masked generation approach [chang2022maskgit]. However, its masking schedule is heuristically determined without a theoretical foundation, and its multi-scale tokenization introduces error accumulation during inference, leading to inferior performance than SOKE. For fairness, we also report the performance of SOKE when using Qwen3-0.6B as the LM, instead of the originally adopted mBART-Large [mbart]. Although both LMs have a similar size (0.6B), we observe that employing Qwen3 leads to slightly better performance, which can be attributed to the larger pretraining corpus of the Qwen3 series.
Furthermore, we report the performance of an MDLM baseline obtained by directly training an MDLM on sign language datasets, without any of our proposed techniques (pretraining, UTC, or MoP). Notably, this baseline performs on par with SOKE, which utilizes external sign dictionaries, while offering 40% higher throughout (Fig.˜6a). This highlights the potential of MDLMs for SLG. Finally, our proposed MaDiS achieves new SOTA results across the test sets of all three datasets and all evaluation metrics.
Figure 6: (a) Comparison of throughput measured in output tokens per second. (b-d) Efficiency-quality trade-off on three datasets. The generation length is fixed to 100 tokens, and the number of sampling steps is set to 100, 50, 25 (default), and 10, corresponding to decoding 1, 2, 4, and 10 tokens per step, respectively.
Figure 7: (a) User study with professional CSL signers. We report average ratings of the previous SOTA method, SOKE [zuo2025soke], our proposed MaDiS, ground-truth motions, and ground-truth RGB videos. Error bars indicate 95% confidence intervals. (b) Absolute values of Spearman’s rank correlation coefficient [spearman] between signers’ ratings and metric scores. DTW and SiBLEU refer to the average of Body and Hands. (c) Visualization of attention maps from the penultimate Transformer layer.
Efficiency. As shown in Fig.˜6a, across all test samples of the three datasets, our method achieves a throughput of 132 tokens/sec, representing a 40% improvement over SOKE with KV caching. This high throughput enables the generation of a 400-frame (100-token) video in only 0.76s, thereby satisfying real-time requirements. This improvement stems from the highly efficient parallel token generation enabled by the reverse process in MDLMs.
In Fig.˜6b-d, we observe that MaDiS provides a flexible trade-off between generation quality and efficiency, where quality is measured by DTW and SiBLEU scores (averaged over Body and Hands). When the number of sampling steps is set to 100, corresponding to sampling one token at a time, the model achieves the highest generation quality but the lowest throughput. In contrast, with only 10 sampling steps, inference becomes extremely fast (exceeding 300 tokens/sec) at the cost of degraded performance. Using 25 sampling steps yields a balance between generation quality and efficiency.
Human Evaluation. To assess the intelligibility of the generated signs, we conducted a user study involving five professional CSL signers with 4 to 13 years of signing experience. They were asked to rate the alignment between generated signs and text annotations on a scale from 1 to 10, where higher scores indicate higher sign accuracy and better semantic conveyance. Ground-truth motions and RGB videos were also included for reference. Specifically, we provided 15 generated signs from SOKE and from our proposed MaDiS, together with the corresponding ground-truth motions and videos. The order of all samples was randomly shuffled to avoid potential bias. All participants were compensated above the average hourly rate in China.
As shown in Fig.˜7a, our method achieves an average rating of 6.9, substantially outperforming SOKE, which receives an average rating of 4.6. The ground-truth motions obtain an average rating of 8.4, further demonstrating the feasibility of using motions as effective sign language representations. Furthermore, we report the absolute values of Spearman’s rank correlation coefficient [spearman] (1 means high correlation while 0 means no correlation) to study the alignment of the metric scores with human preference (signers’ ratings). As shown in Fig.˜7b, SiBLEU and SiCLIP exhibit strong correlation, whereas DTW, based on minimum-distance alignment, shows weaker correlation.